The energy required to remove both electrons from the Helium from its ground state is 79eV. Ho much energy is required to ionize Helium(remove one electron)?
1
Expert's answer
2020-06-02T17:51:37-0400

The energy required to remove one electron from the ground state of "He^+" (Hydrogen like atom )atom is

"E_n=\frac{13.6 Z^2}{n^2}" eV

For ground state "n=1" ,

"E_1=\frac{13.6 (2)^2}{1^2}=54.4 eV"

Given that the energy required to remove both electrons from the Helium from its ground state is "79eV" .

So the energy required to ionize Helium is

= "79 eV -54.4 eV=24.6 eV"